OpenStudy (4815162342):

How many liters of 15.0 molar NaOH stock solution will be needed to make 20.0 liters of a 1.2 molar NaOH solution? Show the work used to solve this problem.

OpenStudy (aveline):

\[\frac{ 20.0L }{ 1 }\times \frac{ 1.2 moles NaOH }{ L.ofsolution }\times \frac{ L.ofsolution }{ 15.0 moles NaOH }\]

OpenStudy (aveline):

How many liters will it take?

OpenStudy (4815162342):

1.6? @Aveline

OpenStudy (aveline):

Correct